OpenStudy (anonymous):

Evaluate the following expression using the given values: Find x - 3y if x = 3 and y = -2. Please help :(

OpenStudy (babalooo):

what's wrong? just replace x with 3 and y with -2 in the equation

OpenStudy (anonymous):

This is very easy all you do is put the numbers in where they go

OpenStudy (anonymous):

3-3(-2)= whats 3x(-2) i have no clue how do you multiply a negative and a positive

OpenStudy (amistre64):

multiplication means we add a value a certain number of times ... not that this is a neg times a pos mind you.

OpenStudy (amistre64):

3(-2) = -2-2-2

OpenStudy (anonymous):

3 isn't exponent its just a number in front

OpenStudy (amistre64):

an exponent means that we multiply a value a certain number of times ....

OpenStudy (babalooo):

note the "ADD" a certain value number of times

OpenStudy (anonymous):

-6? is that what 3x(-2)?

OpenStudy (amistre64):

3x(-2) = -2-2-2 = -6 yes

OpenStudy (anonymous):



OpenStudy (amistre64):

tinker, you made an error

OpenStudy (anonymous):

huh?

OpenStudy (amistre64):

site keeps giving me a 'lost connection' memo ...

OpenStudy (babalooo):

me too :(

OpenStudy (babalooo):

see how it is 3-[(3)(-2)] you found what was in the brackets which was -6 3-(-6) becomes 3+6

OpenStudy (anonymous):

Ok so what did i do wrong @amistre64

OpenStudy (amistre64):

3 - (-6) not 3 - 6

OpenStudy (anonymous):

Ooh double negative...I forgot that, thank you :)

OpenStudy (ericanoel912):

its 3-(-6), so 3+6= 9

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (ericanoel912):

no problem :)